OpenStudy (anonymous):

Someone help please I'll give a medal.One solution of the equation 6x^2 − 19x + 3 = 0 is 1/6. What is the other solution?

OpenStudy (anonymous):

Help please best answer gets a medal!

OpenStudy (anonymous):

cant figure it out?

OpenStudy (anonymous):

solve 6x^2 − 19x + 3 = 0

OpenStudy (anonymous):

Type that in into http://www.wolframalpha.com/

OpenStudy (anonymous):

Great answer to find solutions and figure out how to get the answer

OpenStudy (anonymous):

does it show work?

OpenStudy (anonymous):

Not for these problems

OpenStudy (anonymous):

Here is still a great place to get explanations But do you know how to use quadratic formula?

OpenStudy (anonymous):

Oh okay ill see if I can figure it out. Thank you for your help

OpenStudy (ja1):

@m_sequence1 This is a study website where we help users learn, and learn ourselves, please do not post websites like those, if you can you can guide the user to the answer and explain, but please do not post answering websites.

OpenStudy (anonymous):

Alright sounds good

OpenStudy (ja1):

0 = 6x^2 − 19x + 3 Find a GCF, is there one?

OpenStudy (anonymous):

yes 3?

OpenStudy (anonymous):

\[\div2a\]

OpenStudy (anonymous):

$$6x^2 − 19x + 3 = 0\\x^2-\frac{19}6+\frac12=0$$Note since we have a second-degree polynomial we expect \(2\) roots and therefore solutions to the above equation; recall that our constant is the product of our roots, so if we let \(c\) be our unknown root we have:$$\frac12=\frac16c\\c=3$$

OpenStudy (ja1):

Yes good, so now what do we get?

OpenStudy (anonymous):

dam my equations didn't come out

OpenStudy (anonymous):

\(-\frac{19}6x\)

OpenStudy (anonymous):

thanks everyone for your help I wish i could give multiple medals lol ill give to oldrin.bataku but oldrin could you give it to JA1 thanks guys

OpenStudy (ja1):

No need @Wowzers :) he did all the work, I just corrected @m_sequence1

OpenStudy (phi):

when they give you a solution you know x= 1/6 multiply both sides by 6 6x = 1 add -1 to both sides 6x -1 = 0 this means (6x-1) is one of the factors divide 6x-1 into the original quadratic to find the other factor

OpenStudy (anonymous):

thats phi for help but I got it :)